The undetermined coefficient set of y-2y+9y–18y = 2* +x?is a) {e2+,x?,x,1} b) {xe*,x,x1} c) {x?e2*, **, x,1} d) {xe2\, x,
Answer #1

The undetermined coefficient set of y—24+gy–189 = é 4x? solution To find the undeterinined coefficients we have to find th2u de IF put D=2 2x ue I = (3076D+9) 으리 ve (12-879) 13 Calculating Ig GR - Ig +9 وم) (2-0) 2 x (0+3) Ig = 2 X (it D 57 - (D Dap² Ig - 2 81 azh co 2 dre I q 20 e = 20 e I 28 124 ) 81 - 9 I Integrating الاوے In = f* [62ceno) C142)+}{re) I= -) 4 a so 1,